# The $S_{n}$-equivariant Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$

### v1

@@ Abstract.
We compute the $S_{n}$-equivariant topological Euler characteristic of the Kontsevich moduli space ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$. Letting ¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)\subset\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ denote the subspace of maps from curves without rational tails, we solve for the motive of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ in terms of ¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)$ and plethysm with a genus-zero contribution determined by Getzler and Pandharipande [GP06]. Fixing a generic $\mathbb{C}^{\star}$-action on $\mathbb{P}^{r}$, we derive a closed formula for the Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ as an $S_{n}$-equivariant virtual mixed Hodge structure, which leads to our main formula for the Euler characteristic of ¯$\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$. Our approach connects the geometry of torus actions on Kontsevich moduli spaces [GP99] with symmetric functions in Coxeter types $A$ and $B$ [Mac95], as well as the enumeration of graph colourings with prescribed symmetry [Han85]. We also prove a structural result about the $S_{n}$-equivariant Euler characteristic of ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ in arbitrary genus.
@@ 1. Introduction
Let ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ denote the Kontsevich moduli space of stable maps of degree $d$ from $n$-pointed curves of genus $g$ to $\mathbb{P}^{r}=\mathbb{P}^{r}_{\mathbb{C}}$. It is a fundamental object in enumerative geometry and in the study of algebraic curves. The fibres of the map ¯$\overline{\mathcal{M}}_{g}(\mathbb{P}^{r},d)\to\overline{\mathcal{M}}_{g}$ are studied in Brill–Noether theory. While the moduli space is smooth and irreducible when $g=0$, it is usually reducible, non-equidimensional, and highly singular when $g>0$. These pathologies make it difficult to study the geometry of ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ in positive genus. From the enumerative perspective, these issues have been circumvented via torus localisation by Graber and Pandharipande [GP99], following earlier work of Kontsevich [Kon95]. They express the Gromov–Witten invariants of $\mathbb{P}^{r}$ in terms of computable integrals on the space ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ of torus-fixed stable maps. The action of $\mathbb{C}^{\star}$ is induced by a generic one on $\mathbb{P}^{r}$, with $r+1$ isolated fixed points.
In this paper, we enhance this perspective, shifting from integrals to $S_{n}$-equivariant motivic invariants of the moduli space ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$, where the symmetric group $S_{n}$ permutes the marked points. This space admits a combinatorial stratification, where each stratum is isomorphic to a product of moduli spaces of stable curves, modulo the action of the automorphism group of a finite graph. In the case $g=1$, we solve the combinatorics of the strata, at the motivic level, in terms of plethysm of symmetric functions. Our main application is a formula for the $S_{n}$-equivariant topological Euler characteristic
@@
be the generating function for the $S_{n}$-equivariant Euler characteristics. When $g=0$, a recursive formula for $\mathsf{b}_{0,r}$ has been given by Getzler and Pandharipande [GP06]. As noted in §1.5 below, their result is actually strong enough to determine the $S_{n}$-equivariant Hodge decomposition of the moduli space. In positive genus, the pathologies of the space make it a subtle problem to calculate the coarser invariant $\mathsf{b}_{g,r}$.
We determine $\mathsf{b}_{1,r}$ in terms of $\mathsf{b}_{0,r}^{\prime}$, as well as the two terms $\mathsf{a}_{0},\mathsf{a}_{1}\in\Lambda$, where
@@
Closed formulas for $\mathsf{a}_{g}$ when $g\leq 1$ have been given by Getzler: see [Get95a, Theorem 5.7] for $\mathsf{a}_{0}$ and [Get99, Corollary 5.4] for $\mathsf{a}_{1}$. The terms $\mathsf{a}_{0}$ and $\mathsf{a}_{1}$ are combined in our formula to determine $\mathsf{b}_{1,r}^{\mathrm{nrt}}$, which is defined by
@@
where ¯$\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)\subset\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ is the subspace of maps from smooth elliptic curves or nodal cycles of smooth rational curves. Such curves are characterised as the ones which do not contain any rational tails, as the superscript indicates. To state our main theorem, we recall that $\Lambda=\mathbb{Q}[[p_{1},p_{2},\ldots]],$ where $p_{n}$ is the $n$th power sum symmetric function. For a graded symmetric function $f\in\Lambda[[q]]$ we put $f^{\prime}:=\frac{\partial f}{\partial p_{1}}$ and $\dot{f}:=\frac{\partial f}{\partial p_{2}}$. We use the symbol $\circ$ to denote plethysm of graded symmetric functions. For $f\in\Lambda[[q]]$, we set $\psi_{n}(f):=p_{n}\circ f$. We use the notation $\varphi$ for Euler’s totient function and $\mu$ for the Möbius function on the integers.
@@
$+\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\eta_{k,d}(r)\frac{(1+2\dot{\mathsf{a}}_{0})^{2}}{(1-\psi_{2}(\mathsf{a}_{0}^{\prime\prime}))^{k/2+1}}+\sum_{j\mid k}\theta_{j,k,d}(r)\frac{1}{(1-\psi_{j}(\mathsf{a}_{0}^{\prime\prime}))^{k/j}}.$
@@
$\eta_{k,d}(r)=\frac{1}{4}\sum_{j\mid k}\sum_{i\mid j}\sum_{\begin{subarray}{c}{\ell}\\ {2\ell\mid i\text{ and }k\mid d\ell}\end{subarray}}\mu\left(\frac{i}{2\ell}\right)\binom{\frac{d\ell}{k}-1}{\ell-1}(r^{\ell+1}+r^{\ell}),$
@@
Theorem A determines the generating function $\mathsf{b}_{1,r}$ after substituting the known formulas for $\mathsf{a}_{0}$, $\mathsf{a}_{1}$, and $\mathsf{b}_{0,r}^{\prime}$ discussed above. Sample calculations obtained using Theorem A are included in Tables 1, 2, and 3 in §7. For fixed $d$ and $n$, the multiplicity of each irreducible character of $S_{n}$ in $\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d))$ is a polynomial function of $r$, of degree at most $d+1$, with rational coefficients. We generalise our techniques to establish that this structural result holds for ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ in every genus $g$, as stated in Theorem C.
The formula for $\mathsf{b}_{1,r}^{\mathrm{nrt}}$ in Theorem A follows from our calculation of the $S_{n}$-equivariant Serre characteristic [^1] Our notation follows that of [GP06]. The Serre characteristic is a refinement of what has been called the Hodge–Deligne polynomial or $E$-polynomial. of the torus fixed locus
@@
We use tools from both algebraic geometry and algebraic combinatorics to study ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$. We define the key tools and outline how they are applied in more detail below.
@@ 1.1. Graded $\mathbb{S}$-spaces and plethysm
An $\mathbb{S}$-space $\mathcal{X}$ is the data of an $S_{n}$-variety $\mathcal{X}(n)$ for each integer $n\geq 0$, while a graded $\mathbb{S}$-space is an $S_{n}$-variety $\mathcal{X}(n,d)$ for each pair of integers $n,d\geq 0$. The Grothendieck ring $K_{0}(\mathsf{V},\mathbb{S})$ of $\mathbb{S}$-spaces and the Grothendieck ring $K_{0}(\mathsf{V},\mathbb{S})[[q]]$ of graded $\mathbb{S}$-spaces are introduced in [GP06]. The Serre characteristic
@@
[…]
¯$\overline{\mathcal{M}}_{g}^{\mathrm{nrt}}(r)\subset\overline{\mathcal{M}}_{g}(r).$
@@
[¯$[\overline{\mathcal{M}}_{g}(r)]=[\overline{\mathcal{M}}_{g}^{\mathrm{nrt}}(r)]\circ\left([\operatorname{Spec}\mathbb{C}]+\frac{[D\overline{\mathcal{M}}_{0}(r)]}{[\mathbb{P}^{r}]}\right),$
@@
Since a recursive formula for $\mathsf{e}(\overline{\mathcal{M}}_{0}(r))$ has been determined in [GP06], this formula reduces the calculation of $\mathsf{e}(\overline{\mathcal{M}}_{g}(r))$ to that of $\mathsf{e}(\overline{\mathcal{M}}_{g}^{\mathrm{nrt}}(r))$.
[…]
@@
for the Serre characteristic of the locus of $\mathbb{C}^{\star}$-fixed genus one maps without rational tails. The starting point for our calculation of $\mathscr{B}_{r}$ is the graph stratification of ¯$\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ described by Graber and Pandharipande [GP99]. The strata correspond to $(\mathbb{P}^{r})^{\mathbb{C}^{\star}}$-coloured, edge-weighted graphs of total weight $d$, which we call localisation graphs. If $\mathcal{G}$ is a localisation graph, then the action of $\operatorname{Aut}(\mathcal{G})$ on the set of half-edges $H(\mathcal{G})$ of $\mathcal{G}$ induces an embedding
@@
[…]
@@
[¯$[\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)^{\mathbb{C}^{\star}}]=[(\mathbb{P}^{r})^{\mathbb{C}^{\star}}]\cdot[\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(0)]+[\mathsf{Dih}_{r}]\circ_{S_{2}}[\mathsf{Cat}].$
@@
The term ¯$\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(0)$ parameterises maps to a point, and its Serre characteristic has been computed [Get98, Pet12], as has the Serre characteristic of $\mathsf{Cat}$ [BM13]. Theorem 4.1 thus reduces the calculation of $\mathscr{B}_{r}$ to the calculation of $\mathsf{e}(\mathsf{Dih}_{r})$ and its wreath product plethysm with $\mathsf{e}(\mathsf{Cat})$. The Serre characteristic $\mathsf{e}(\mathsf{Dih}_{r})$ is valued in ${K_{0}(\mathsf{M})\otimes\Lambda(S_{2})[[q]]}$, where $\Lambda(S_{2})$ is the ring of wreath product symmetric functions studied by Macdonald [Mac95].

### v2

# The $S_{n}$-equivariant Euler characteristic of $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$

@@ Abstract.
We compute the $S_{n}$-equivariant topological Euler characteristic of the Kontsevich moduli space $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$. Letting $\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)\subset\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ denote the subspace of maps from curves without rational tails, we solve for the motive of $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ in terms of $\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)$ and plethysm with a genus-zero contribution determined by Getzler and Pandharipande [25]. Fixing a generic $\mathbb{C}^{\star}$-action on $\mathbb{P}^{r}$, we derive a closed formula for the Euler characteristic of $\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ as an $S_{n}$-equivariant virtual mixed Hodge structure, which leads to our main formula for the Euler characteristic of $\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$. Our approach connects the geometry of torus actions on Kontsevich moduli spaces [31] with symmetric functions in Coxeter types $A$ and $B$ [43], as well as the enumeration of graph colourings with prescribed symmetry [33].
@@ 1. Introduction
Let $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ denote the Kontsevich moduli space of stable maps of degree $d$ from $n$-pointed curves of genus $g$ to $\mathbb{P}^{r}=\mathbb{P}^{r}_{\mathbb{C}}$. It is a fundamental object in enumerative geometry and in the study of algebraic curves. The fibres of the map $\overline{\mathcal{M}}_{g}(\mathbb{P}^{r},d)\to\overline{\mathcal{M}}_{g}$ are studied in Brill–Noether theory. While the moduli space is smooth and irreducible when $g=0$, it is usually reducible, non-equidimensional, and highly singular when $g>0$. These pathologies make it difficult to study the geometry of $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$ in positive genus. From the enumerative perspective, these issues have been circumvented via torus localisation by Graber and Pandharipande [31], following earlier work of Kontsevich [41]. They express the Gromov–Witten invariants of $\mathbb{P}^{r}$ in terms of computable integrals on the space $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ of torus-fixed stable maps. The action of $\mathbb{C}^{\star}$ is induced by a generic one on $\mathbb{P}^{r}$, with $r+1$ isolated fixed points.
In this paper, we enhance this perspective, shifting from integrals to $S_{n}$-equivariant motivic invariants of the moduli space $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$, where the symmetric group $S_{n}$ permutes the marked points. This space admits a combinatorial stratification, where each stratum is isomorphic to a product of moduli spaces of stable curves, modulo the action of the automorphism group of a finite graph. In the case $g=1$, we solve the combinatorics of the strata, at the motivic level, in terms of plethysm of symmetric functions. Our main application is a formula for the $S_{n}$-equivariant topological Euler characteristic
@@
be the generating function for the $S_{n}$-equivariant Euler characteristics. When $g=0$, a recursive formula for $\mathsf{b}_{0,r}$ has been given by Getzler and Pandharipande [25]. As noted in §1.4 below, their result is actually strong enough to determine the $S_{n}$-equivariant Hodge decomposition of the moduli space. In positive genus, the pathologies of the space make it a subtle problem to calculate the coarser invariant $\mathsf{b}_{g,r}$.
We determine $\mathsf{b}_{1,r}$ in terms of $\mathsf{b}_{0,r}$, as well as the two terms $\mathsf{a}_{0},\mathsf{a}_{1}\in\Lambda$, where
@@
Closed formulas for $\mathsf{a}_{g}$ when $g\leq 1$ have been given by Getzler: see [22, Theorem 5.7] for $\mathsf{a}_{0}$ and [24, Corollary 5.4] for $\mathsf{a}_{1}$. The terms $\mathsf{a}_{0}$ and $\mathsf{a}_{1}$ are combined in our formula to determine $\mathsf{b}_{1,r}^{\mathrm{nrt}}$, which is defined by
@@
where $\overline{\mathcal{M}}_{1,n}^{\mathrm{nrt}}(\mathbb{P}^{r},d)\subset\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d)$ is the subspace of maps from smooth elliptic curves or nodal cycles of smooth rational curves. Such curves are characterised as the ones which do not contain any rational tails, as the superscript indicates. To state our main theorem, we recall that $\Lambda=\mathbb{Q}[[p_{1},p_{2},\ldots]],$ where $p_{n}$ is the $n$th power sum symmetric function. For a graded symmetric function $f\in\Lambda[[q]]$ we put $f^{\prime}:=\frac{\partial f}{\partial p_{1}}$ and $\dot{f}:=\frac{\partial f}{\partial p_{2}}$. We use the symbol $\circ$ to denote plethysm of graded symmetric functions. For $f\in\Lambda[[q]]$, we set $\psi_{n}(f):=p_{n}\circ f$. We use the notation $\varphi$ for Euler’s totient function and $\mu$ for the Möbius function on the integers.
@@
$+\sum_{d\geq 2}q^{d}\sum_{k=2}^{d}\left(\eta_{k,d}(r)\frac{(1+2\dot{\mathsf{a}}_{0})^{2}}{(1-\psi_{2}(\mathsf{a}_{0}^{\prime\prime}))^{k/2+1}}+\sum_{j\mid k}\theta_{j,k,d}(r)\frac{1}{(1-\psi_{j}(\mathsf{a}_{0}^{\prime\prime}))^{k/j}}\right).$
@@
$\eta_{k,d}(r)=\frac{1}{4}\sum_{j\mid k}\sum_{\begin{subarray}{c}{\ell}\\ {2\ell\mid j\text{ and }k\mid d\ell}\end{subarray}}\mu\left(\frac{j}{2\ell}\right)\binom{\frac{d\ell}{k}-1}{\ell-1}(r^{\ell+1}+r^{\ell}),$
@@
Theorem A determines the generating function $\mathsf{b}_{1,r}$ after substituting the known formulas for $\mathsf{a}_{0}$, $\mathsf{a}_{1}$, and $\mathsf{b}_{0,r}^{\prime}$ discussed above. Sample calculations obtained using Theorem A are included in Tables 1, 2, and 3 in §7. For fixed $d$ and $n$, the multiplicity of each irreducible character of $S_{n}$ in $\chi^{S_{n}}(\overline{\mathcal{M}}_{1,n}(\mathbb{P}^{r},d))$ is a polynomial function of $r$, of degree at most $d+1$, with rational coefficients. In [36, Theorem C], we show further that this holds for $\chi^{S_{n}}(\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d))$ for any $g$.
The formula for $\mathsf{b}_{1,r}^{\mathrm{nrt}}$ in Theorem A follows from our calculation of the $S_{n}$-equivariant Serre characteristic [^1] Our notation follows that of [25]. The Serre characteristic is a refinement of what has been called the Hodge–Deligne polynomial or $E$-polynomial. of the torus fixed locus
@@
We use tools from both algebraic geometry and algebraic combinatorics to study $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$. We define the key tools and outline how they are applied in more detail below.
@@ 1.1. Graded $\mathbb{S}$-spaces and plethysm
An $\mathbb{S}$-space $\mathcal{X}$ is the data of an $S_{n}$-variety $\mathcal{X}(n)$ for each integer $n\geq 0$, while a graded $\mathbb{S}$-space is an $S_{n}$-variety $\mathcal{X}(n,d)$ for each pair of integers $n,d\geq 0$. The Grothendieck ring $K_{0}(\mathsf{V},\mathbb{S})$ of $\mathbb{S}$-spaces and the Grothendieck ring $K_{0}(\mathsf{V},\mathbb{S})[[q]]$ of graded $\mathbb{S}$-spaces are introduced in [25]. The Serre characteristic
@@
where the composition on the right is plethysm of symmetric functions. Write $\overline{\mathcal{M}}_{g}(r)$ for the graded $\mathbb{S}$-space whose $(n,d)$-component is $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)$. A genus-$g$ stable map $f:C\to\mathbb{P}^{r}$ is said to have no rational tail if the domain curve $C$ does not contain any non-trivial subcurve of arithmetic genus $g$. This is a combinatorial condition on dual graphs and defines a locally closed $\mathbb{S}$-subspace
$\overline{\mathcal{M}}_{g}^{\mathrm{nrt}}(r)\subset\overline{\mathcal{M}}_{g}(r).$
@@
$\overline{\mathcal{M}}_{g}(r)]=[\overline{\mathcal{M}}_{g}^{\mathrm{nrt}}(r)]\circ\left([\operatorname{Spec}\mathbb{C}]+\frac{[D\overline{\mathcal{M}}_{0}(r)]}{[\mathbb{P}^{r}]}\right),$
@@
Since a recursive formula for $\mathsf{e}(\overline{\mathcal{M}}_{0}(r))$ has been determined in [25], this formula reduces the calculation of $\mathsf{e}(\overline{\mathcal{M}}_{g}(r))$ to that of $\mathsf{e}(\overline{\mathcal{M}}_{g}^{\mathrm{nrt}}(r))$.
1.2. $\mathbb{C}^{\star}$-localisation and type-$B$ symmetric functions
In genus one, we express the geometry of the space of torus-fixed maps from curves without rational tails using type-$B$ combinatorics. The full graded $\mathbb{S}$-space $\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)$ has subtle geometry: it includes the space of maps $\mathcal{M}_{1,n}(\mathbb{P}^{r},d)$ from smooth elliptic curves for all $n$ and $d$, as well as maps to $\mathbb{P}^{r}$ from arbitrary cycles of smooth pointed rational curves. The picture simplifies upon passing to the $\mathbb{C}^{\star}$-fixed locus. We set
@@ 1.2. $\mathbb{C}^{\star}$-localisation and type-$B$ symmetric functions
for the Serre characteristic of the locus of $\mathbb{C}^{\star}$-fixed genus one maps without rational tails. The starting point for our calculation of $\mathscr{B}_{r}$ is the graph stratification of $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{r},d)^{\mathbb{C}^{\star}}$ described by Graber and Pandharipande [31]. The strata correspond to $(\mathbb{P}^{r})^{\mathbb{C}^{\star}}$-coloured, edge-weighted graphs of total weight $d$, which we call localisation graphs. If $\mathcal{G}$ is a localisation graph, then the action of $\operatorname{Aut}(\mathcal{G})$ on the set of half-edges $H(\mathcal{G})$ of $\mathcal{G}$ induces an embedding
@@
If we restrict to $\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)$, the graph $\mathcal{G}$ must be a cycle of length $k\geq 2$. Thus (1.6) factors through the hyperoctahedral group $B_{k}:={S_{2}\wr S_{k}}$. Therefore, when specialised to $\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)^{\mathbb{C}^{\star}}$, equation (1.6) assigns to each localisation graph $\mathcal{G}$ a character of $B_{k}$, which encodes the action of $\operatorname{Aut}(\mathcal{G})$ on $H(\mathcal{G})$. We find that the $S_{n}$-action on the stratum determined by $\mathcal{G}$ depends only on this character. We package localisation graphs and their characters into a graded $\mathbb{B}$-space $\mathsf{Dih}_{r}$, where $\mathbb{B}$-spaces are the natural type-$B$ analogues of the $\mathbb{S}$-spaces discussed in §1.1. The moduli in the stratum defined by $\mathcal{G}$ comes from the choice of a pointed chain of rational curves for each vertex. The moduli space of chains is an $S_{2}\times\mathbb{S}$-space $\mathsf{Cat}$, where $S_{2}$ acts by switching the orientation of the chain. We prove the following motivic formula for the graded $\mathbb{S}$-space $\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)^{\mathbb{C}^{\star}}$.
@@
$\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(r)^{\mathbb{C}^{\star}}]=[(\mathbb{P}^{r})^{\mathbb{C}^{\star}}]\cdot[\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(0)]+[\mathsf{Dih}_{r}]\circ_{S_{2}}[\mathsf{Cat}].$
@@
The term $\overline{\mathcal{M}}_{1}^{\mathrm{nrt}}(0)$ parameterises maps to a point, and its Serre characteristic has been computed [23, 54], as has the Serre characteristic of $\mathsf{Cat}$ [7]. Theorem 4.1 thus reduces the calculation of $\mathscr{B}_{r}$ to the calculation of $\mathsf{e}(\mathsf{Dih}_{r})$ and its wreath product plethysm with $\mathsf{e}(\mathsf{Cat})$. The Serre characteristic $\mathsf{e}(\mathsf{Dih}_{r})$ is valued in $K_{0}(\mathsf{M})\otimes\Lambda(S_{2})[[q]]$}, where $\Lambda(S_{2})$ is the ring of wreath product symmetric functions studied by Macdonald [43].